# RoboCerebra: A Large-scale Benchmark for Long-horizon Robotic Manipulation Evaluation

### v1

@@ 2 Related Work
Vision-Language-Action (VLA) Models. Recent advances in VLMs Qwen-VL ; liu2023llava ; reid2024gemini ; openai2024gpt4o ; han2024videoespresso ; liu2024llavanext ; han2023llms and large-scale manipulation datasets bu2025agibot ; o2024open have driven the development of VLA models kim2024openvla ; brohan2022rt that translate language instructions into executable robotic actions. While these models have shown strong performance in short-horizon tasks, their precision is often constrained by discrete action representations. To address this, recent works black2024pi_0 ; hou2025dita ; chi2023diffusion propose using diffusion models to represent continuous action spaces, improving expressiveness and control fidelity. Nonetheless, existing approaches remain limited in their ability to generalize to long-horizon scenarios. Most focus on low-level policy execution or short-term planning, and struggle to maintain coherence across extended temporal contexts. In this work, we adopt OpenVLA as the low-level controller, and focus on how high-level reasoning via hierarchical System 2 planning can enhance performance in long-horizon manipulation tasks.
@@ 3.2 Dataset construction pipeline
Human Demonstration and Annotation. Despite extensive automation, high-quality demonstrations remain essential for learning grounded, long-horizon manipulation. We employ human operators to perform task instructions in simulation, generating diverse and realistic action trajectories. Each trajectory is annotated with fine-grained subtask boundaries to align actions with specific task steps. This subtask-level annotation enables precise temporal segmentation, supports diverse execution styles, and ensures comprehensive coverage across task variants and lengths. It further facilitates grounding of language instructions to motion segments, enabling the training of temporally-aware and instruction-conditioned policies. To maintain annotation quality at scale, we allocate significant human effort: 400 hours for trajectory and time annotation, alongside 200 hours dedicated to a separate verification phase, as outlined in Tab. 2. This rigorous process ensures the consistency, completeness, and accuracy of labeled plans and state transitions.
@@ 5.3 Ablation on Planner
Table 5: Evaluation of System 2 capabilities from multiple perspectives, including Average Planning Accuracy ($Acc_{P}$), Observation Judgment ($Acc_{C}$), Success Rate ($SR$), Average Plan Length ($Len$) and Plan Efficiency ($\eta$).
Model $Acc_{P}$uparrow $Acc_{C}$uparrow $SR$uparrow $Len$downarrow $\eta$uparrow
@@ 5.4 Evaluation on System 2
6 Results on Memory Tasks
@@
VLM $SR_{Exp.}$uparrow $SR_{Exp.-only}$uparrow $\eta_{Exp.}$uparrow $SR_{Exe.}$uparrow $Acc_{Dec.}$uparrow
@@
7 Conclusion
Appendix A Detailed Task Suite
Figure 5: Example from RoboCerebra for different tasks.
Appendix B Prompt Details of Task Generation
@@
Appendix C Prompt Details of Hierarchical Framework
@@
Appendix D Case Study on Planning
@@
Appendix E Case Study on Sub-Plan
@@
Appendix F Case Study on Memory Tasks

### v2

@@ 2 Related Work
Vision-Language-Action (VLA) Models. Recent advances in VLMs Qwen-VL ; liu2023llava ; reid2024gemini ; openai2024gpt4o ; han2024videoespresso ; liu2024llavanext ; han2023llms ; shu2025semantics and large-scale manipulation datasets bu2025agibot ; o2024open have driven the development of VLA models kim2024openvla ; brohan2022rt that translate language instructions into executable robotic actions. While these models have shown strong performance in short-horizon tasks, their precision is often constrained by discrete action representations. To address this, recent works black2024pi_0 ; hou2025dita ; chi2023diffusion propose using diffusion models to represent continuous action spaces, improving expressiveness and control fidelity. Nonetheless, existing approaches remain limited in their ability to generalize to long-horizon scenarios. Most focus on low-level policy execution or short-term planning, and struggle to maintain coherence across extended temporal contexts. In this work, we adopt OpenVLA as the low-level controller, and focus on how high-level reasoning via hierarchical System 2 planning can enhance performance in long-horizon manipulation tasks.
@@ 3.2 Dataset construction pipeline
Human Demonstration and Annotation. Despite extensive automation, high-quality demonstrations remain essential for learning grounded, long-horizon manipulation. We employ human operators to perform task instructions in simulation, generating diverse and realistic action trajectories. Each trajectory is annotated with fine-grained subtask boundaries to align actions with specific task steps. This subtask-level annotation enables precise temporal segmentation, supports diverse execution styles, and ensures comprehensive coverage across task variants and lengths. It further facilitates grounding of language instructions to motion segments, enabling the training of temporally-aware and instruction-conditioned policies. To maintain annotation quality at scale, we allocate significant human effort: 400 hours for trajectory and time annotation, alongside 200 hours dedicated to a separate verification phase, as outlined in Tab. 3.2. This rigorous process ensures the consistency, completeness, and accuracy of labeled plans and state transitions.
@@ 5.3 Ablation on Planner
Table 5: Evaluation of System 2 capabilities from multiple perspectives, including Average Planning Accuracy ($Acc_{P}$), Observation Judgment ($Acc_{C}$), Success Rate ($SR$), Average Plan Length (LenLen) and Plan Efficiency ($\eta$).
Model $Acc_{P}$ $\uparrow$ $Acc_{C}$ $\uparrow$ SRSR $\uparrow$ LenLen $\downarrow$ $\eta$ $\uparrow$
@@ 5.4 Evaluation on System 2
6 Conclusion
7 Acknowledgements
This research is supported in part by National Key R&D Program of China (2022ZD0115502), Ningbo Science and Technology Innovation 2025 Major Project (2025Z034), National Natural Science Foundation of China (NO. 62461160308, U23B2010), "Pioneer" and "Leading Goose" R&D Program of Zhejiang (No. 2024C01161). In addition, this research was supported in part by NUS Start-up Grant A-0010106-00-00.
Appendix A Appendix
A.1 Detailed Task Suite
Figure 5: Example from RoboCerebra for different tasks.
A.2 Study on Different VLA models
Table 6: Performance comparison of VLMs under OpenVLA and $\pi$0-fast. Results show that the current VLM maintains comparable performance across both settings, indicating that its System 2 capability is independent of the specific VLA architecture.
Method Avg Ran. Obs. Exp. Exe. Mix Ideal
Qwen2.5-LM-OpenVLA 11.87 18.90 12.88 7.02 10.87 2.55 18.90
Qwen2.5-LM-$\pi$0-fast 11.47 16.32 11.94 9.92 8.27 4.73 17.63
LLaVA-N-Blind-OpenVLA 8.00 13.97 12.33 3.54 3.54 0.37 14.25
LLaVA-N-Blind-$\pi$0-fast 5.99 7.89 8.06 5.89 3.85 1.82 8.42
GPT-4o-Blind-OpenVLA 15.10 20.00 17.03 7.02 16.09 10.48 20.00
GPT-4o-Blind-$\pi$0-fast 13.63 14.47 15.82 12.27 12.69 11.27 15.26
Qwen2.5-VL-OpenVLA 11.19 14.25 14.25 2.63 12.61 6.67 16.71
Qwen2.5-VL-$\pi$0-fast 13.19 20.79 15.45 8.24 6.15 6.91 21.58
LLaVA-N-video-OpenVLA 11.37 16.71 16.16 1.07 10.87 3.70 19.73
LLaVA-N-video-$\pi$0-fast 8.79 12.11 12.12 7.73 4.81 2.55 13.42
GPT-4o+OpenVLA 16.04 18.63 19.45 8.04 16.69 11.48 21.92
[…]
GT-plan-OpenVLA 25.16 26.85 30.68 19.47 23.48 19.26 31.23
GT-plan-$\pi$0-fast 23.04 23.68 26.36 18.15 16.92 26.55 26.58
To verify whether the System 2 reasoning capability of the current VLM depends on a specific VLA architecture, we conducted additional experiments comparing two System 1 modules — OpenVLA and $\pi$0-fast (implemented following the openpi repository).
As shown in Table 6, the model exhibits comparable performance under both configurations across all evaluated metrics. This consistency suggests that the higher-level cognitive and planning abilities of the VLM function independently of the underlying low-level control mechanism. Although slight variations can be observed in metrics such as Execution and Exploration, the overall performance trend remains stable. These results demonstrate that the reasoning and decision-making processes of the VLM are robust and transferable across different embodied control implementations, confirming the generality and architecture-independence of its System 2 module.
A.3 Study on Different Planners
Table 7: Comparison of specialized embodied VLMs with general-purpose models. Specialized VLMs show stronger performance on Embodied-QA benchmarks and complex scenarios (Exp. and Mix), though their abilities have not fully generalized to Embodied Planning.
Method Para. Avg Ran. Obs. Exp. Exe. Mix Ideal
LLaVA-N-Blind 7B 8.00 13.97 12.33 3.54 3.54 0.37 14.25
Cosmos-Reason1 7B 8.41 7.63 10.45 5.55 7.31 8.73 10.79
VeBrain 8B 9.41 12.89 12.35 7.06 3.65 4.21 16.32
Qwen2.5-VL 7B 11.19 14.25 14.25 2.63 12.61 6.67 16.71
LLaVA-N-Video 7B 11.37 16.71 16.16 1.07 10.87 3.70 19.73
RoboBrain-2.0 7B 11.40 12.11 12.06 9.92 10.96 7.27 16.05
GPT-4o - 16.04 18.63 19.45 8.04 16.69 11.48 21.92
As shown in Table 7, VLMs specifically designed for embodied tasks—such as RoboBrain-2.0, Cosmos-Reason1, and VeBrain—generally outperform general-purpose VLMs on the Embodied-QA benchmark. These specialized models also exhibit superior adaptability in our constructed complex tasks, particularly in challenging scenarios such as Memory Exploration and Mix, which require long-horizon reasoning and dynamic decision-making.
However, this specialization has not yet fully transferred to the domain of Embodied Planning, where limitations remain in generating coherent planning sequences and decomposing actions into fine-grained steps. This observation suggests that, while task-specific VLMs demonstrate stronger contextual understanding and perception-grounded reasoning, their higher-level planning and general reasoning abilities are still developing. These results highlight an important direction for future research on embodied intelligence — achieving deeper integration between perception, reasoning, and long-term planning.
A.4 Results on Memory Tasks
@@
VLM $SR_{Exp.}$ $\uparrow$ $SR_{Exp.-only}$ $\uparrow$ $\eta_{Exp.}$ $\uparrow$ $SR_{Exe.}$ $\uparrow$ $Acc_{Dec.}$ $\uparrow$
@@
A.5 Prompt Details of Task Generation
@@
A.6 Prompt Details of Hierarchical Framework
@@
A.7 Case Study on Planning
@@
A.8 Case Study on Sub-Plan
@@
A.9 Case Study on Memory Tasks